OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3

OpenStudy (anonymous):

Its easily.. take the 2nd eq and solve it for y and then plug that value in eq 1 and solve for x..

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

x-y=3 x=3+y plug in eq 1. 4(3+y)+2y=6 12+4y+2y=6 6y=6-12 y=-1 plug it in eq 1 and solve for x: 4x+2(-1)=6 x=2